OpenStudy (anonymous):

How many H atoms are present in a 7.98ng sample of caffeine C8H10O2N4?

OpenStudy (anonymous):

\[7.98 ng(\frac{1 g}{1x10^{-9}ng})(\frac{1 mole comp}{194g comp})(\frac{6.022x10^{23}molec}{1 mole comp})(\frac{10 H atoms}{1 molec})=?\]

OpenStudy (anonymous):

put that in the calculator and you should get your answer. Also check the molar mass of the compound I googled it because i am lazy.

OpenStudy (osanseviero):

Zbay can you help me after this please?

OpenStudy (anonymous):

Why did you divide 7.98 by 1x10^-9? Shouldn't you multiply it to get grams?

OpenStudy (anonymous):

yes flip that

OpenStudy (anonymous):

it should be \[\frac{1x10^{-9}g}{1 ng}\]

OpenStudy (anonymous):

okay.. thank you so much! :)